OpenStudy (anonymous):

can someone explain to me how to simplify expression and write as positive exponent (2x/y)^-3

OpenStudy (chaise):

2xy^3 Recall that 1/x=x^(-1) Also vice versa. 1/x^(-1)=x^(1)=x

OpenStudy (chaise):

Alternatively, 1/x^(-3)=x^3 You can apply this to your problem to give 2xy^3